Solve:
5x/7 – 6x/5 = 3/35

Answer by checkley71(8403)

5X/7-6X/5=3/35 FIND A COMMON DENOMINATOR FOR THE LEFT SIDE FRACTIONS THUS (7*5=35)
(5X*5-6X*7)/35=3/35
(25X-42X)/35=3/35
-7X/35=3/35
-7X=3
X=3/-7
X=-3/7